A recipe for trail mix calls for 1 cup of raisins for every 2 cups of granola. Write an equation that describes the relationship between raisins and granola. Then graph the linr
1
Expert's answer
2020-12-20T18:51:37-0500

Let "x=" the number of cups of granola and "y=" the number of cups of raisins.

At the beginning we have "x=0" cups of granola and "y=0" cups of raisins.

We have the direct proportion "y=mx, x\geq0."

Given that a recipe calls for "y=1" cup of raisins for every "x=2" cups of granola


"1=m(2)=>m=\dfrac{1}{2}"

The equation that describes the relationship between raisins "(y)" and granola "(x)" is


"y=\dfrac{1}{2}x, x\geq0"